# Advancements in Functorial Homological Mirror Symmetry

### v3

@@ Abstract
Mostly inspired by recent work by Katzarkov, Kontsevich, and Sheshmani, combined with previous work by Aganagic, Ooguri, Saulina and Vafa with regard to BPS black hole microstate counting in terms of topological field theory calculations, we will show how these tools can be applied to concrete setups arising from String Theory, and why the formalism of functorial Homological Mirror Symmetry needs to be further developed.
@@ 1 Introduction and motivation
Derived and noncommutative geometry is a paradigm in which to formulate and study various dualities pioneered by String Theory, including Mirror Symmetry. In the context of topological string theory, the relevant physical theory is an extended 2D toplogical field theory (TFT). A central problem in noncommutative geometry is that of extracting the partition function of the TFT on various manifolds from the data of the category, and to describe various geometric operations on the tangent spaces for the A- and B-model TFT in purely categorical terms. Relying on some recent advancements in this topic, connecting them to calculations in perturbative topological String Theory we highlight possible directions for further extending the functorial formulation of Homological Mirror Symmetry.
The present work is structured as follows: Section 2 introduces symmetric obstruction theories as shifted symplectic structures, DO ; PTVV . Section 3 briefly overviews a recent work by Katzarkov, Kontsevich and Sheshmani, KKS , pertaining the categorification of DT-invariants for cases involving non-smooth moduli spaces, relative CY structures and spherical functors. In Section 4 we then turn to the topological string theory formulation of Section 3, mostly related to Pan ; AOSV . Section 5 applies the tools outlined in the previous sections to a particular setup of interest for its applicability to String Theory, namely that of Moore-Tachikawa varieties, MT ; MS ; P ; CM . Particular emphasis is placed on the role of Coulomb branches for classifying 2D TFTs, XY ; CT ; CT1 ; BFN . In Section 6 we will then turn to some of the main techniques used for extending HMS for singular varieties. We conclude by outlining currently ongoing work by the same author towards extending the formalism in question.
Throughout the treatment, we highlight several possible directions for further development that we plan to be addressing next.
@@ 2.1 Shifted symplectic and Lagrangian structures
$j^{*}:j^{*}\left(\mathbb{L}_{{}_{F/k}}\right)\ \rightarrow\ \mathbb{L}_{{}_{F/k}}$ (2)
@@
It is an ${\cal O}$-compact object endowed with an ${\cal O}$-orientation of dimension 3. Then, the derived stack of perfect complexes on $X$, $\mathbb{R}$Perf$(X)$, is classically endowed with a (-1)-shifted symplectic structure, thereby defining a symmetric obstruction theory on the truncation
@@ 2.1.1 PTVV
The derived critical locus of any function carries a shifted symplectic structure, since Crit$(f)$ is the intersection of the zero section with the graph of $df$ in the cotangent bundle.
[…]
@@
[…]
@@ 3 Categorification of DT-invariants
We start from briefly overviewing the geometric setup of KKS  [^6] Addressing the question of S-duality arising in Superstring Theory.. Given a smooth projective CY3-variety, $X$. Assuming Pic$(X)$ is generated by an ample divisor $L$. For a fixed integer $k\ \in\ \mathbb{Z}_{{}_{>0}}$, $H\ \in\ |kL|$, we get a sheaf of $H$-modules over the $H$-hypersurface [^7] Where D-branes are placed., whose fiber jumps in dimension upon crossing $H$.
@@
$\text{Ch}^{{}^{\bullet}}(i,n)\ =\ \left(0,[H],\frac{H^{{}^{2}}}{2}-i\ell,\chi(O_{{}_{H}})-H\frac{d_{{}_{2}}(X)}{2}-n\right).$ (31)
@@
In principle, the smoothing features infinite cohomologies. However, we can always identify a map, $\phi$, such that ${\cal A}_{{}_{smooth}}$ is obtained from a variety with at most two cohomologies
$\mathbb{E}^{{}^{\times}}\ \xrightarrow{\phi}\ {\cal A}_{{}_{smooth}}.$ (33)
@@
[…]
$[R]\ =\ \text{number}\ \cdot\ [\text{pt}],$ (36)
@@ 3.1 Tyurin degeneration
[…]
Figure 2: Pictorial representation of how the moduli space splits into two under Tyurin degeneration. The LHS is associated to a CY threefold, $X$, whereas the RHS is associated to the union of two Fano varieties, $Y_{{}_{1}},Y_{{}_{2}}$, joined together by an anticanonical divisor, $S$.
@@
$\text{Pref}\left(Y_{{}_{1}}\right)\underset{\text{Perf}(X)}{\times}\text{Pref}\left(Y_{{}_{2}}\right)$ (44)
to the category Perf$(X)$. The latter being CY, leads one to assume (44) to be CY as well.
@@
=$=\int\prod_{{}_{i}=1}^{{}^{|2g-2|}}dU_{{}_{i}}\left|Z_{{}_{\text{top}}}\left(U_{{}_{1}},...,U_{{}_{|2g-2|}}\right)\right|^{{}^{2}},$
@@ 4.2 TFT formulation of gluing rules
[…]
@@ 5 Cohomological Field Theories (CohFTs)
    We then turn to a particular realisation of quantum cohomology in K-theory, 6.2. The reason being the analogous treatment performed by Moore and Segal, MS , and its applicability to the categorification of D-branes in String Theory. Section 5.2 constitutes a specific application of the formalism initiated by MT , also known as the Moore-Tachikawa varieties.

### v4

@@ Abstract
Mostly inspired by recent work by Katzarkov, Kontsevich, and Sheshmani, combined with previous work by Aganagic, Ooguri, Saulina and Vafa with regard to BPS black hole microstate counting in terms of topological field theory calculations, we will show how these tools can be applied to concrete setups arising from String Theory, and why the formalism of functorial Homological Mirror Symmetry needs to be further developed. A crucial ingredient will turn out being cobordism techniques for evaluating invariants.
@@ 1 Introduction and motivation
Derived and noncommutative geometry is a paradigm in which to formulate and study various dualities pioneered by String Theory, including Mirror Symmetry. In the context of topological string theory, the relevant physical theory is an extended 2D toplogical field theory (TFT). A central problem in noncommutative geometry is that of extracting the partition function of the TFT on various manifolds from the data of the category, and to describe various geometric operations on the tangent spaces for the A- and B-model TFT in purely categorical terms. Relying on some recent advancements in this topic, connecting them to calculations in perturbative topological String Theory, we highlight possible directions for further extending the functorial formulation of Homological Mirror Symmetry.
As we shall see, two main pillars of our analysis rely on the following correspondences, that have been put forward in the past few decades [^2] RW and AG in these diagrams are a shorthand notation for Rozansky-Witten and Algebraic Geometry, respectively.:
$\text{Gauging}\ \ \ \ \ \ \ \longleftrightarrow\ \ \ \ \ \ \ \text{deformation quantisation}$ (1)
$\text{RW-model}\ \ \ \ \ \ \ \longleftrightarrow\ \ \ \ \ \ \ \text{Categorified AG}$ (2)
How such correspondences fit within the purpose of our work will be explained in due course. Our main finding consists in having related stability and transversality in the degeneracy formula for non-ideal sheaves to nonabelian gauging in the RW-theory setup.
The present work is structured as follows: Section 2 introduces symmetric obstruction theories as shifted symplectic structures, DO ; PTVV . Section 3 briefly overviews a recent work by Katzarkov, Kontsevich and Sheshmani, KKS , pertaining the categorification of Donaldson-Thomas (DT) invariants for cases involving non-smooth moduli spaces, relative CY structures and spherical functors. The example we outline falls within the right-hand-side of (2). We highlight the types of mathematical structures we need to apply to achieve the degeneracy formula, and how it relates to deformation quantisation. In Section 4 we then turn to the topological string theory formulation of Section 3, mostly related to Pan ; AOSV . Here we stress the role played by different brane types in the braneweb for which the topological partition function is being calculated, and how they are encoded in a 2-functor corresponding to Gromov-Witten (GW) invariants [^3] DT-invariants are an equivariant formulation of GW-invariants. Calculations coincide for CY3-folds.. Cobordism calculations, though, highlight that such invariants restrict to critical points of a topological Morse function. More interesting configurations should be expected to arise when attempting to analyse geometric structures arising at arbitrary points of the flow. Section 5 applies the tools outlined in the previous sections to a specific setup of interest for its applicability to String Theory, namely that of Moore-Tachikawa varieties, MT ; MS ; P ; CM . Particular emphasis is placed on the role of Coulomb branches for classifying 2D Topological Field Theories (TFTs), XY ; CT ; CT1 ; BFN . The reason for doing so is that such 2D TFTs constitute the boundary of a 3D TFT described by Rozansky and Witten for describing the category of branes associated to the B-model and their deformation. Section 6 turns to a major technique used for extending HMS for singular varieties, namely that of abelianisation. In so doing, we explain the relation between gauging and deformation quantisation, (1), and the difference between the abelian and the nonabelian setup, highlighting how this relates to the definition of the virtual fundamental class in the degeneracy formula. We conclude by outlining some main open questions and currently ongoing work by the same author towards extending the formalism in question.
@@ 2.1 Shifted symplectic and Lagrangian structures
$j^{*}:j^{*}\left(\mathbb{L}_{{}_{F/k}}\right)\ \rightarrow\ \mathbb{L}_{{}_{h^{{}^{0}}(F)/k}}$ (4)
@@
It is an $\cal O$}-compact object endowed with an $\cal O$}-orientation of dimension 3. Then, the derived stack of perfect complexes on $X$, $\mathbb{R}$Perf(X)(X), is classically endowed with a (-1)-shifted symplectic structure, thereby defining a symmetric obstruction theory on the truncation
@@ 2.1.1 PTVV
The derived critical locus of any function carries a shifted symplectic structure, since Crit(f)(f) is the intersection of the zero section with the graph of $df$ in the cotangent bundle.
Theorem 3 [PTVV, 2013] Let $(X,\omega)$ be a derived stack with an $n$-shifted symplectic structure, and let $L\rightarrow X$ and $L^{{}^{\prime}}\rightarrow X$ be morphisms of derived stacks equipped with Lagrangian structures. Then, $L\underset{X}{\times}L^{{}^{\prime}}$ is equipped with a natural (n1)(n-1)-shifted symplectic structure.
@@
$\Lambda_{{}_{1}}\ \times_{{}_{\mathfrak{g}^{*}/G}}\ \Lambda_{{}_{2}}\ =\ 0/G\ \times_{{}_{\mathfrak{g}^{*}/G}}\ M/G\ \simeq\ \mu^{{}^{-1}}(0)/G.$ (26)
@@ 3 Categorification of DT-invariants
[…]
@@
$\text{Ch}^{{}^{\bullet}}(i,n)\ =\ \left(0,[H],\frac{H^{{}^{2}}}{2}-i\ell,\chi(O_{{}_{H}})-H\frac{Tod(X)}{2}-n\right).$ (33)
@@
In principle, the smoothing features infinite cohomologies. However, we can always identify a map, $\phi$, such that ${\cal A}_{{}_{smooth}}$ is obtained from a variety with at most two cohomologies. At the level of the cotangent bundle, it reads as follows
$\mathbb{E}^{{}^{\times}}\ \xrightarrow{\phi}\ \mathbb{A}_{{}_{smooth}}.$ (35)
@@
To the neighbourhood of the deformation [^10] RR actually denotes the set of points that are invariant under deformation., $R$, there is an assigned cycle, $[R]$, which, in turn can be decomposed as follows
[R]=number[pt],[R]\ =\ $\text${number}\ $\cdot\$$\text${pt}], (38)
@@
3.0.1 Hypersurface rigidity: advantages and caveats
The main issue encountered upon evaluating DT-invariants is ensuring transversality and stability are simultaneously satisfied.
Given a fibration
$H\ \hookrightarrow$ $X$ (42)
$\downarrow\ \pi$
$\mathbb{P}^{{}^{1}}$
there is a moduli space of Hodge structures for $X$ and $H$. The latter introduces a relative moduli space of Hodge structures for the fibration
$X$ (43)
$\downarrow$
$\mathbb{P}^{{}^{1}}$
corresponding to
$p\ \in\ \mathbb{P}\ \mapsto\ H^{{}^{2}}\left(X_{{}_{p}},K_{{}_{X_{{}_{p}}}}\right)$ (44)
DT-invariants ultimately count only the curves surviving the deformation. (42) can be extended to the case where $H\equiv K3$, in which case the partition function $Z_{{}_{i}}^{{}^{H}}(q)$ is known to be a modular form. Rigidity of the surface $H$ simplifies a lot the calculation. For example, assuming $X$ to be a quintic 3-fold in $\mathbb{CP}^{{}^{4}}$, one can take $K=1,L=H=\mathcal{O}(1)$ to be an ample divisor. Focusing only on ideal sheaves on $H$, the aim is that of making the information on the hypersurface $H$ as simple as possible. Specifically, one could look at 0-torsion sheaves, which could be realised in two different ways since they share the same Chern characters. Counting only rank-1 contributions, $K=1$. One could further assume the sheaf, $\mathcal{F}$, to be the pushforward of 1D schemes of $H$ in $X$.
Figure 2: The picture on the left shows a moduli space of Hodge structures. Picking a class $\beta\in H^{{}^{2}}(H,\mathbb{Z})$, $D^{{}^{\beta}}$ is a divisor in this analytic structure that encodes the curve (in red). On the right hand side, a curve $C$ (in blue) inside the hypersurface $H$ (in red) is taken as the union of points and a continuous curve.
Take a curve and points on $H$
$C\ =\ Z\ \cup\ P_{{}_{1}}\ \cup...\ \cup\ P_{{}_{n}}\ \subset\ H\ \subset\ X$ (45)
This induces a short-exact sequence
[…]
with $\mathcal{I}_{{}_{C/H}}$ denoting a rank-1 ideal on $H$. The Chern characters on the complement of the curve, restricted to the hypersurface, reads
$\text{Ch}\left(\mathcal{I}_{{}_{C/H}}\right)\ =\ \left(1,-\beta,-n\right)$ (47)
Pushing forward to $X$, (47) reads
$\text{Ch}\left(\iota_{{}_{*}}\mathcal{I}_{{}_{C/H}}\right)\ =\ \left(0,H,-\beta,-n\right)$ (48)
which is rank-0 on the curve, and rank-1 everywhere else. Due to the codimension-2 nature of the curve with respect to $X$, we can construct a commutative diagram between the two ideal sheaves, that further induces the following short-exact sequence
[…]
Importantly, $\left[\mathcal{O}(-H)\ \rightarrow\ \mathcal{I}_{{}_{C/X}}\right]$ and $\iota_{{}_{*}}\mathcal{I}_{{}_{C/X}}$ are quasi-isomorphic to each other in the derived category, implying they deform similarly as long as
$H^{{}^{i}}\left(\mathcal{O}_{{}_{X}},\mathcal{I}_{{}_{C/H}}(H)\right)\ =\ 0\ \ \ \forall\ i>0$ (50)
But these objects are always Joyce-Song stable. Computational invariance implies that, starting from the moduli space, we have the following projective fibration
$M\left(\mathcal{O}_{{}_{X}}(-H)\ \longrightarrow\ \mathcal{I}_{{}_{C/X}}\right)\ \xrightarrow{\mathbb{P}\left(H^{{}^{\bullet}}\left(X,\mathcal{I}_{{}_{C/X}}(H)\right)\right)}\ \text{Hilb}\left(C\subset X\right)\ =\ M\left(\mathcal{I}_{{}_{C/X}}\right)$ (51)
At the level of DT-invariants, (51) leads to factorisation as follows
$\text{DT}\left(M\left(\mathcal{O}_{{}_{X}}(-H)\ \longrightarrow\ \mathcal{I}_{{}_{C/X}}\right)\right)\ =\ \text{DT}\left(M\left(\mathcal{I}_{{}_{C/X}}\right)\right)\cdot\ \chi\left(\mathbb{P}\left(H^{{}^{\bullet}}\left(X,\mathcal{I}_{{}_{C/X}}(H)\right)\right)\right)$ (52)
But this is assuming rigidity of $H$. In particular, rigidity in this setup leads to failure in evaluating all terms in the expansion of $Z_{{}_{i}}^{{}^{H}}(q)$. Furthermore, this method features shortcomings if one were to account for deformation of subschemes.
A further setup that one could be considering is for $H=K3$, analysed by [Ch., Sheshmani], requiring assuming Serre duality in the deformation obstruction theory. The main caveat is that allowing the surface to deform, the ideal sheaf of points stays the same, but the Cartier divisor might disappear. This is relevant in Hodge Theory, in the setting that given cohomology/homology, can we uniquely specify the corresponding variety? There might be no holomorphic cycles supporting certain curves. We therefore need to ensure that a certain deformation preserves all curves (associated to the D2-branes), not just the points (D0-branes). Need to count fibers in (43), where $\beta$ after deformation can be represented by a holomorphic curve. Counting such things defines the Noether-Lefschetz number ($NL$)
$Z_{{}_{i}}^{{}^{H}}(q)=\ Z\left(\text{Hilb}^{{}^{n}}(H)\right)\cdot\ NL$ (53)
With such construction, KKS proved the S-modularity conjecture for the case of a compact CY, a surface, and sheaves supported on the surface. However, general CY3 complete intersections or quintic 3-folds, the study of S-duality requires using the degeneration technique, corresponding to taking a variety, degenerating it as follows
$X_{{}_{5}}\ \ \ \longrightarrow\ \ \ X_{{}_{4}}\ \cup\ \mathbb{P}^{{}^{3}}$ (54)
where $X_{{}_{5}}$ is the zero-locus of a function $f_{{}_{5}}$,
$X_{{}_{5}}\ \overset{def.}{=}\ Z\left(f_{{}_{5}}\left(x_{{}_{0}},...,x_{{}_{4}}\right)\right)$ (55)
Then, one can construct a family that is degeneration-equal to the original setup (55)
$tf_{{}_{5}}\left(x_{{}_{0}},...,x_{{}_{4}}\right)\ -\ f_{{}_{1}}\left(x_{{}_{0}},...,x_{{}_{4}}\right)\ +\ f_{{}_{2}}\left(x_{{}_{0}},...,x_{{}_{4}}\right)\ \ \ \ \forall t$ (56)
implying
$Z\left(f_{{}_{1}}\right)\ =\ \mathbb{P}^{{}^{3}}\ \ \ \ \ \ ,\ \ \ \ \ \ Z\left(f_{{}_{2}}\right)\ =\ X_{{}_{4}}$ (57)
Consistency with (54) is readily proved by showing that, for $t=0$, (56) reduces to a product of a quartic and a hyperplane, whereas for $t\neq 0$ we get the original variety, $X_{{}_{5}}$.
(54) can therefore be rewritten as two Fano varieties meeting at an anticanonical divisor under Tyurin degeneration
$X_{{}_{5}}\ \ \ \longrightarrow\ \ \ X_{{}_{4}}\ \underset{K3}{\cup}\ \mathbb{P}^{{}^{3}}$ (58)
with more specific details outlined in Section 3.1. However, for the formalism to apply, we need the space of degenerating family to be smooth, which is not due to the presence of a base locus. To smooth the family, one needs to perform a blow-up procedure, which is a common technique in birational geometry
$X_{{}_{5}}\ \ \ \longrightarrow\ \ \ X_{{}_{4}}\ \cup\ B\underset{S}{|}\mathbb{P}^{{}^{3}}$ (59)
@@ 3.1 Tyurin degeneration
$Z:\ X\ \rightarrow\ Y_{{}_{1}}\ \underset{S}{\bigcup}\ Y_{{}_{2}}.$ (63)
Figure 3: Pictorial representation of how the variety splits into two under Tyurin degeneration. The LHS is associated to a CY threefold, $X$, whereas the RHS is associated to the union of two Fano varieties, $Y_{{}_{1}},Y_{{}_{2}}$, joined together by an anticanonical divisor, $S$.
@@
$\text{Perf}\left(Y_{{}_{1}}\right)\underset{\text{Perf}(X)}{\times}\text{Perf}\left(Y_{{}_{2}}\right)$ (64)
to the category Perf(X)(X). The latter being CY, leads one to assume (64) to be CY as well.
@@
$=\int\prod_{{}_{i}=1}^{{}^{|2g-2|}}dU_{{}_{i}}\left|Z_{{}_{\text{top}}}\left(U_{{}_{1}},...,U_{{}_{|2g-2|}}\right)\right|^{{}^{2}},$
@@
3.1.1 Sheave splitting
Any sheaf $\mathcal{F}\in X$ or $\mathcal{F}_{{}_{i}}\in Y_{{}_{i}}$, $i=1,2\in\text{Coh}(\mathbb{P})$. There exists a short-exact-sequence
[…]
This is where the issue arises: tensoring a codimension-1 sheaf with something that is also codimension-1. This is not the usual tensor product, in particular, it does not lie in $D^{{}^{b}}\text{Coh}(X)$. From this, it thereby follows that the $\mathcal{F}_{{}_{i}}$ sheaves are not necessarily meeting $S$ transversally, in particular the relative setting features a nontrivial torsion element
$\text{Tor}_{{}_{1}}^{{}^{\mathcal{O}_{{}_{Y_{{}_{i}}}}}}\left(\mathcal{F}_{{}_{i}},\mathcal{O}_{{}_{S}}\right)\ \neq\ 0$ (69)
Requiring compactification to make sheaves transverse
$M\left(Y_{{}_{i}}/S,\text{Ch}_{{}_{i}}^{{}^{*}}\right)\ \Subset\ M\left(Y_{{}_{i}},\text{Ch}_{{}_{i}}^{{}^{*}}\right)$ (70)
Alternatively to the blow-up, one can perform acordionical expansion, then gluing. Correspondingly, the curve (45) splits into two parts (with respect to the relative moduli spaces)
$M(C\hookrightarrow X)\ \rightarrow\ \overline{M\left(C_{{}_{1}}\hookrightarrow Y_{{}_{1}}/S\right)}\ \cup\ \overline{M\left(C_{{}_{2}}\hookrightarrow Y_{{}_{2}}/S\right)}$ (71)
with respective virtual classes
$\left[\bar{M}\left(C_{{}_{1}}\hookrightarrow Y_{{}_{1}}/S\right)\right]\ \ \ \ \ \ ,\ \ \ \ \ \ \left[\bar{M}\left(C_{{}_{2}}\hookrightarrow Y_{{}_{2}}/S\right)\right]$ (72)
$\left[M_{{}_{1}}\underset{\text{Hilb}^{{}^{n}}}{\times}M_{{}_{2}}\right]^{{}^{\text{vir}}}\ \overset{def.}{=}\ \Delta^{{}^{!}}\left(\left[M_{{}_{1}}\right]^{{}^{\text{vir}}}\times\left[M_{{}_{2}}\right]^{{}^{\text{vir}}}\right)\ \equiv\ \left[M\left(C\hookrightarrow\ X\right)\right]^{{}^{\text{vir}}}$ (73)
with the last equality following by deformation invariance.
When it is fully transversal, the Hilbert polynomial factorises
$\text{Hilb}^{{}^{n_{{}_{1}}}}(S)\times\text{Hilb}^{{}^{n_{{}_{2}}}}(S)$ (74)
Upon further relaxing the assumption of the sheaves being ideal, the degeneracy technique fails in its ordinary setup, and one therefore needs to resort to derived intersection theory, instead, KKS . Considering the moduli space for DG-schemes of rigidified perfect complexes, one can achieve a suitably upgraded counterpart to (58), on which the shifted symplectic structure formalism can still be applied as explained in Section 2.
For the purpose of our work, the main feature of the result by KKS is that the degeneracy formula in presence of non-ideal sheaves, can be achieved and relates two DG-categories that are related by deformation quantisation. However, as will be explained in Section 6, this is not yet the most general setup that one could consider. Specifically, it still falls within the abelianised setup.
@@ 4.1 Black holes and CY $X=L_{{}_{1}}\oplus L_{{}_{2}}\rightarrow\Sigma_{{}_{g}}$
The D4-D2-D0-brane system is a configuration that can be realised by a suitable choice of a curve $C$ on $H\subset X$. In Section 3, we highlighted the importance of assuming such curve to encode all the information of the CY, $X$, to calculate the corresponding DT-invariant. Likewise, the same thing can be stated for the calculation of the topological partition function, (91), where all the information is assumed to be contained in the specified brane system.
@@ 4.2 TFT formulation of gluing rules
$\int dU\ \text{Tr}_{R_{{}_{1}}}U\ \text{Tr}_{R_{{}_{1}}}U^{{}^{-1}}\ =\ \delta_{{}_{R_{{}_{1}}R_{{}_{2}}}}.$ (101)
Figure 4 corresponds to critical points of the systole function. But, in principle, we should explore beyond the critical points, namely considering considering configurations at arbitrary points of the flow, interpolating between different SOTs.
@@ 5 Cohomological Field Theories (CohFTs)
    We then turn to a particular realisation of quantum cohomology in K-theory, C. The reason being the analogous treatment performed by Moore and Segal, MS , and its applicability to the categorification of D-branes in String Theory. Section 5.2 constitutes a specific application of the formalism initiated by MT , also known as the Moore-Tachikawa varieties.